# Consumption Partial Insurance

### v4

@@ 2 Lifecycle model with tail income risk
where $U(\cdot;\mathbf{Z}_{it})$ is utility over consumption, $\mathbf{Z}_{it}$ is a vector of observed taste shifters (e.g., age, education), $\beta$ is the discount factor, $r$ is the deterministic interest rate, $A_{i0}$ is beginning-of-time financial wealth, ${\operatorname*{\mathbb{E}}}_{0}$ denotes expectations over uncertain future income $Y_{it}$, and $T$ is the end of the horizon. We assume that utility satisfies standard regularity conditions, in particular it is continuously twice differentiable, but we do not otherwise parameterize it.
@@ 2.1 Income process
where $\mathbf{X}_{it}$ is a vector of characteristics in the $t=0$ information set of the household that drive the deterministic profile of income over the lifecycle (e.g., age, education, race); $\boldsymbol{\delta}$ is the loading factor of those variables on the logarithm of income. The remaining terms make up the stochastic component of income, consisting of permanent income $P_{it}$ and transitory shock $v_{it}$. We assume that permanent income has a unit root, $P_{it}=P_{it-1}+\zeta_{it}$, where $\zeta_{it}$ is the permanent income shock. These modeling choices give rise to the familiar permanent-transitory formulation
$\Delta y_{it}=\zeta_{it}+\Delta v_{it},$ (3)
@@
We allow the central moments of shocks to depend on $t$ to reflect that different stages of the lifecycle may exhibit different amounts of dispersion, skewness, or thickness of the tails of income risk; Guvenen et al. (2021) provide some evidence for that. Otherwise, shocks across households are identically distributed given $t$. Permanent and transitory shocks are mutually independent and independent over time.
@@ 2.2 Linear consumption function
where $\Delta c_{it}$ is the residual from a regression of real consumption growth, $\Delta\ln C_{it}$, on the taste observables $\mathbf{Z}_{it}$ (i.e., consumption growth net of its predictable profile). The linear function (4) stems from a log-linearization (i.e., a first-order Taylor series expansion) of the problem’s first-order conditions and the household budget constraint, as we show in appendix A. These approximations are independent of the distributions of income shocks and mimic similar derivations in BPP (who abstract from higher-order moments).
@@
The transmission parameters vary, in principle, with $i$ (household) and $t$ (age), reflecting heterogeneity in financial wealth, among other things. [^8] In addition, households in Blundell et al. (2016) differ in their access to external insurance and insurance through family labor supply, while households in Theloudis (2021) differ in their preferences. The idea is that a household can use wealth to smooth consumption when income shocks hit; and the more wealth one holds the better they can self-insure. We establish this link from wealth to $\phi_{it}^{(1)}$, $\psi_{it}^{(1)}$ in appendix A. So while the pass-through of shocks depends on wealth, it does not depend on the magnitude or sign of the shock itself. Function (4) thus implies that a household can insure big or small, good or bad shocks similarly, a point to which we return subsequently.
@@ 2.3 Quadratic consumption function
The linear consumption function offers an approximation to the true consumption response to income shocks when such shocks do not move the household too far away from its deterministic consumption profile. Function (4) relies on a linearization of marginal utility (the problem’s first-order conditions) around that deterministic profile; but such a linearization will be inaccurate in the presence of tail income risk, e.g., when extreme shocks may induce large shifts in consumption as a consequence of concave utility. [^9] This echoes Carroll (2001)’s critique of the log-linearized Euler equation. Kaplan and Violante (2010) and Blundell et al. (2013) assess the approximations in BPP across alternative scenarios, though not in one with respect to the severity of income shocks.
@@
where $\Delta c_{it}$ and $\xi_{it}$ are defined as previously. The quadratic function (5) stems from a second-order Taylor approximation to the problem’s first-order conditions and the household budget constraint, as we show in appendix A. These operations and resulting expressions generalize the original approximations in BPP and in several papers thereafter. [^10] We do not go beyond a quadratic approximation for reasons of tractability and because higher-order terms in a regression of $\Delta c_{it}$ on a polynomial in $\Delta y_{it}$ are imprecisely estimated.
@@
$\frac{\partial\Delta c_{it}}{\partial\zeta_{it}}=\phi_{it}^{(1)}+2\phi_{it}^{(2)}\zeta_{it}+\omega_{it}^{(22)}v_{it},$ (6)
which depends on multiple parameters and on the magnitude and sign of the income shocks. As such, there is an asymmetric pass-through of good versus bad shocks into consumption. Suppose, for example, that $\phi^{(1)}>0$ and $\phi^{(2)}<0$, as our results subsequently suggest. At $v_{it}=0$ (the average transitory shock), a negative permanent shock has a higher pass-through to consumption relative to a positive shock of the same magnitude. The pass-through of negative shocks would thus be underestimated (i.e., excessive insurance concluded) if the econometrician assessed the extent of partial insurance on the basis of $\phi^{(1)}$ alone, ignoring the non-linearity in the consumption response. The size of this bias increases with the magnitude of $\phi^{(2)}$ and the size of the permanent shock itself. When income risk is non-Gaussian (e.g., it is left tailed), the realization of shocks can be far from their mean, thereby causing large errors in the inference of the degree of partial insurance based on $\phi^{(1)}$ alone.
@@
$\frac{\partial\Delta c_{it}}{\partial v_{it}}=\psi_{it}^{(1)}+2\psi_{it}^{(2)}v_{it}+\omega_{it}^{(22)}\zeta_{it},$ (7)
@@ 3 Identification
Our specification of the income and consumption processes imposes restrictions on the second and higher-order moments of income and consumption that can be used to identify the various parameters. The set of parameters includes the variance ($\sigma_{\zeta_{t}}^{2}$, $\sigma_{v_{t}}^{2}$), skewness ($\gamma_{\zeta_{t}}$, $\gamma_{v_{t}}$), and kurtosis ($\kappa_{\zeta_{t}}$, $\kappa_{v_{t}}$) of the distributions of income shocks, [^12] With a slight abuse of terminology, we use the terms skewness and kurtosis to refer to the third and fourth central moments. We will be clear whenever those terms refer to standardized moments. the partial insurance parameters in the linear ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$) and quadratic consumption functions ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$, $\phi^{(2)}_{t}$, $\psi^{(2)}_{t}$, $\omega^{(22)}_{t}$), and the moments of the unobserved taste heterogeneity ($\sigma_{\xi_{t}}^{2}$, $\gamma_{\xi_{t}}$, $\kappa_{\xi_{t}}$). We initially restrict the partial insurance parameters to not vary across households but we subsequently explore their variation across certain demographic groups and by the household’s position in the distribution of income.
For identification, we assume existence of up to fourth moments of the joint distribution of income and consumption growth. In the case of the quadratic consumption function, we also assume that income shocks are not Bernoulli. For consistent estimation of such moments, particularly over demographic groups subsequently, we further assume non-vanishing shares in each group asymptotically and knowledge of group membership. To carry out statistical inference, we finally assume existence of up to eighth moments.
@@ 3.1 Income process parameters
$\kappa_{v_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta y_{it})^{2}\times(\Delta y_{it+1})^{2})-\sigma_{\zeta_{t}}^{2}\sum_{\tau=0}^{1}\sigma_{v_{t+\tau}}^{2}-\sigma_{\zeta_{t+1}}^{2}\sum_{\tau=-1}^{0}\sigma_{v_{t+\tau}}^{2}-\sigma_{v_{t-1}}^{2}\sum_{\tau=0}^{1}\sigma_{v_{t+\tau}}^{2}$
$-\sigma_{\zeta_{t}}^{2}\sigma_{\zeta_{t+1}}^{2}-\sigma_{v_{t}}^{2}\sigma_{v_{t+1}}^{2}.$
@@ 3.2 Linear consumption function parameters
The moments of the distribution of consumption growth reflect the underlying distributions of income shocks, as well as the distribution of unobserved taste heterogeneity. In general, larger dispersion, skewness, or kurtosis in the distributions of income shocks implies larger dispersion, skewness, or kurtosis in the distribution of consumption growth. The link between the two depends on the degree of insurance with respect to these shocks, so these moments provide restrictions that help identify the parameters of partial insurance.
BPP discuss identification in the linear function at length; we provide a summary here while we relegate some details to appendix B. Formally, the transmission parameter of permanent shocks is identified as
@@
in both cases, the variance of shocks is given by the earlier expressions. [^13] The variance of consumption taste heterogeneity is further given by $\sigma_{\xi_{t}}^{2}={\operatorname*{\mathbb{E}}}((\Delta c_{it})^{2})-(\phi_{t}^{(1)})^{2}\sigma_{\zeta_{t}}^{2}-(\psi_{t}^{(1)})^{2}\sigma_{v_{t}}^{2}$. Its skewness and kurtosis are similarly identified from the higher moment of consumption growth. These statements reflect the intuition behind identification in BPP: the covariance between consumption and (appropriately defined) income growth captures the strength of the relation between the two series and, consequently, the pass-through of income shocks into consumption. [^14] An equivalent view is that the partial insurance parameters are identified from a regression of consumption growth on income growth, using permanent or future income as instruments for income growth.
@@ 3.3 Quadratic consumption function parameters
As per the quadratic specification in (5) and the properties of shocks, the second moment of the distribution of consumption growth is given by
@@
where square matrix $\mathbf{A}$ depends exclusively on the second, third, and fourth moments of income shocks. [^15] The matrix of coefficients is given by $\mathbf{A}=\left(\begin{array}[]{lllll}\sigma_{\zeta_{t}}^{2}&\sigma_{v_{t}}^{2}&\gamma_{\zeta_{t}}&\gamma_{v_{t}}&0\\ \gamma_{\zeta_{t}}&\gamma_{v_{t}}&\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{v_{t-1}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t-1}}^{2})&2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ 0&-\sigma_{v_{t}}^{2}&0&-\gamma_{v_{t}}&0\\ 0&\gamma_{v_{t}}&\sigma_{\zeta_{t}}^{2}(\sigma_{\zeta_{t+1}}^{2}+\sigma_{v_{t}}^{2}+\sigma_{v_{t+1}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(\sigma_{\zeta_{t+1}}^{2}+\sigma_{v_{t+1}}^{2})&0\\ 0&-\gamma_{v_{t}}&-\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}&-\kappa_{v_{t}}&-\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ \end{array}\right).$ , [^16] The variance of consumption taste heterogeneity is identified from ${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{2})$. Its skewness and kurtosis enter the third and fourth moments of consumption growth. Those depend on higher than fourth-order moments of income, which we do not model, so they are generally not identified in the current setting. We establish in appendix B that matrix $\mathbf{A}$ is full rank and invertible for any distribution of income shocks except Bernoulli (for which there is no empirical evidence anyway). [^17] If the distributions of both shocks are Bernoulli, then there exist some parameter configurations for which matrix $\mathbf{A}$ is not full rank. Assuming away Bernoulli distributions is thus an identifying assumption. Identification of the parameters does not require those shocks to be non-Gaussian; the matrix is invertible even if the distributions of shocks are normal and have the same variance between them (i.e., $\sigma_{\zeta_{t}}^{2}=\sigma_{v_{t}}^{2}$).
@@
In the nonlinear specification, tail shocks are associated with a different pass-through from that of average shocks, regardless of the precise nature of the underlying distribution of shocks. The transmission parameters, however, can no longer be identified with second-order income moments alone – inspect matrix $\mathbf{A}$ in (8). Skewness and kurtosis, both defined also in the Gaussian case, are crucial for identifying the pass-through of such shocks, which is intuitive: we want to identify the impact of tail shocks, so we must use information on said tails. In appendix B, we investigate what different levels of income skewness and kurtosis imply for the transmission parameters.
@@
Contrary to income, no higher-order moments of consumption are needed; the identifying moments in (8) are covariances between consumption growth and various powers of income growth. This is an advantage over quantile methods that require observing the entire consumption distribution, e.g., ABB, because, while we may estimate higher moments of consumption growth in the PSID, there is no established population benchmark to assess them against. This is in contrast to higher income moments, for which Guvenen et al. (2021) is a natural benchmark given their use of administrative population data.
@@
This fails to identify the true pass-through of permanent shocks in the presence of nonlinear transmission, i.e., expression (6). Moreover, this workhorse covariance is biased for $\phi_{t}^{(1)}$ if permanent shocks are non-Gaussian. If $\gamma_{\zeta_{t}}<0$, i.e., permanent shocks are left skewed as we find subsequently, and $\phi_{t}^{(2)}<0$, as our empirical estimates suggest, linearity biases $\phi_{t}^{(1)}$ upwards. Similarly, the identifying condition for $\psi_{t}^{(1)}$ in BPP becomes
@@
Our estimation procedure, which we describe in the next section, is linked transparently to our identification argument. Effectively, we bring (8) to the data, along with some over-identifying restrictions, using moments estimation.

### v5

@@ 2 An organizing framework: a lifecycle model with tail income risk
where $U(\cdot;\mathbf{Z}_{it})$ is utility over consumption, $\mathbf{Z}_{it}$ is a vector of observed taste shifters (e.g., age, education), $\beta$ is the discount factor, $r$ is the deterministic interest rate, $A_{i0}$ is initial financial wealth, ${\operatorname*{\mathbb{E}}}_{0}$ denotes expectations over uncertain future income $Y_{it}$, and $T$ is the end of the horizon. We assume that utility satisfies standard regularity conditions, in particular it is continuously twice differentiable, but we do not otherwise parameterize it.
@@ 2.1 Income process
where $\mathbf{X}_{it}$ is a vector of characteristics in the $t=0$ information set of the household that drive the deterministic profile of income over the lifecycle (e.g., age, education, race) and $\boldsymbol{\delta}$ is its loading factor. The remaining terms make up the stochastic component of income, consisting of permanent income $P_{it}$ and transitory shock $v_{it}$. We assume that permanent income has a unit root, $P_{it}=P_{it-1}+\zeta_{it}$, where $\zeta_{it}$ is the permanent income shock. These modeling choices give rise to the familiar permanent-transitory formulation
$\Delta y_{it}=\zeta_{it}+\Delta v_{it},\vspace{-1ex}$ (3)
@@
We allow the central moments of shocks to depend on $t$ to reflect that different stages of the lifecycle may differ in the amount and nature of income risk; Guvenen et al. (2021) provide some evidence for that. Otherwise, shocks across households are identically distributed given $t$. Permanent and transitory shocks are mutually independent and independent over time.
@@ 2.2 Linear consumption function
where $\Delta c_{it}$ is the residual from a regression of real consumption growth, $\Delta\ln C_{it}$, on the taste observables $\mathbf{Z}_{it}$ (i.e., consumption growth net of its predictable profile). The linear function (4) is consistent with a log-linearization (i.e., a first-order Taylor series expansion) of the problem’s first-order conditions and the household budget constraint, as we show in appendix A. These approximations are independent of the distributions of income shocks and mimic similar derivations in BPP (who abstract from higher-order moments).
@@
The transmission parameters vary, in principle, with $i$ (household) and $t$ (age), reflecting heterogeneity in financial wealth, among other things. [^9] In addition, households in Blundell et al. (2016) differ in their access to external insurance and insurance through family labor supply, while households in Theloudis (2021) differ in their preferences. The idea is that a household can use wealth to smooth consumption when income shocks hit; and the more wealth one holds, the better they can self-insure. We establish this link from wealth to $\phi_{it}^{(1)}$, $\psi_{it}^{(1)}$ in appendix A. So while the pass-through of shocks depends on wealth, it does not depend on the magnitude or sign of the shock itself. Function (4) thus implies that a household can insure big or small, good or bad shocks similarly, a point to which we return subsequently.
@@ 2.3 Quadratic consumption function
The linear function offers an approximation to the true consumption rule when income shocks do not move the household too far away from its deterministic consumption profile. Function (4) is consistent with a linearization of marginal utility around that deterministic profile; but such a linearization will be inaccurate in the presence of tail income risk, e.g., when large shocks may induce large shifts in consumption as a consequence of concave utility. [^10] This echoes Carroll (2001)’s critique of approximate Euler equations. He argues that the -then- popular log-linear Euler equation, and to a lesser extent its quadratic counterpart, imply consumption behavior that deviates from that in the true underlying model. However, Kaplan and Violante (2010) and Blundell et al. (2013) assess the linear approximations in BPP across alternative scenarios, though not in one with respect to the severity of income shocks, and they find that they perform reasonably well. The second-order approximations herein aim at improving the inference drawn from these previously established approximations.
@@
where $\Delta c_{it}$ and $\xi_{it}$ are defined as previously. The quadratic function (5) is consistent with a second-order Taylor approximation to the problem’s first-order conditions and the household budget constraint, as we show in appendix A. These operations and resulting expressions generalize the original approximations in BPP and in several papers thereafter. [^11] We do not go beyond a quadratic approximation for reasons of tractability and because higher-order terms in a regression of $\Delta c_{it}$ on a polynomial in $\Delta y_{it}$ are imprecisely estimated.
@@
$\partial\Delta c_{it}/\partial\zeta_{it}=\phi_{it}^{(1)}+2\phi_{it}^{(2)}\zeta_{it}+\omega_{it}^{(22)}v_{it},$ (6)
which depends on multiple parameters and on the magnitude and sign of the income shocks. As such, there is an asymmetric pass-through of good versus bad shocks into consumption. Suppose, for example, that $\phi^{(1)}>0$ and $\phi^{(2)}<0$, as our results subsequently suggest. At $v_{it}=0$ (the average transitory shock), a negative permanent shock has a larger pass-through to consumption relative to a positive shock of the same magnitude. The pass-through of negative shocks would thus be underestimated (i.e., excessive insurance concluded) if the econometrician assessed the extent of insurance on the basis of $\phi^{(1)}$ alone, ignoring the non-linearity in the consumption response. The size of this bias increases with the magnitude of $\phi^{(2)}$ and the size of the negative permanent shock (e.g., the extent of left skewness).
@@
$\partial\Delta c_{it}/\partial v_{it}=\psi_{it}^{(1)}+2\psi_{it}^{(2)}v_{it}+\omega_{it}^{(22)}\zeta_{it},$ (7)
@@
A final remark is needed here. The specific lifecycle model (1)-(2) is not the only structure that gives rise to the quadratic specification (5) or the subsequent econometric framework. Several models, e.g., models with liquidity constraints (Campbell and Hercowitz, 2019) or models for the wealthy hand-to-mouth (Kaplan and Violante, 2014), to offer just two examples, imply consumption policy rules of the form $C_{it}=g_{t}(Y_{it},\dots)$, motivating our empirical specification through a quadratic approximation to $g_{t}$. This light dependence on specific theory is why we used the term ‘quasi-reduced form’ to name our framework.
@@ 3 Identification
Our specification of the income and consumption processes imposes restrictions on the second and higher-order moments of income and consumption that can be used to identify the various parameters. These include the variance ($\sigma_{\zeta_{t}}^{2}$, $\sigma_{v_{t}}^{2}$), skewness ($\gamma_{\zeta_{t}}$, $\gamma_{v_{t}}$), and kurtosis ($\kappa_{\zeta_{t}}$, $\kappa_{v_{t}}$) of the distributions of income shocks, [^13] With a slight abuse of terminology, we use the terms skewness and kurtosis to refer to the third and fourth central moments. We will be clear whenever those terms refer to standardized moments. the partial insurance parameters in the linear ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$) and quadratic consumption functions ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$, $\phi^{(2)}_{t}$, $\psi^{(2)}_{t}$, $\omega^{(22)}_{t}$), and the moments of unobserved taste heterogeneity ($\sigma_{\xi_{t}}^{2}$, $\gamma_{\xi_{t}}$, $\kappa_{\xi_{t}}$). We first restrict the partial insurance parameters to not vary across households but we subsequently explore their variation across certain demographic groups and by the household’s position in the income distribution.
For identification, we assume existence of up to fourth moments of the joint distribution of income and consumption growth. In the case of the quadratic consumption function, we also assume that income shocks are not Bernoulli. For consistent estimation of such moments, particularly over demographic groups subsequently, we further assume non-vanishing shares in each group asymptotically. For inference, we assume existence of up to eighth moments.
@@ 3.1 Income process parameters
$\kappa_{v_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta y_{it})^{2}\times(\Delta y_{it+1})^{2})-\sigma_{\zeta_{t}}^{2}\sum_{\tau=0}^{1}\sigma_{v_{t+\tau}}^{2}-\sigma_{\zeta_{t+1}}^{2}\sum_{\tau=-1}^{0}\sigma_{v_{t+\tau}}^{2}-\sigma_{v_{t-1}}^{2}\sum_{\tau=0}^{1}\sigma_{v_{t+\tau}}^{2}-\sigma_{\zeta_{t}}^{2}\sigma_{\zeta_{t+1}}^{2}-\sigma_{v_{t}}^{2}\sigma_{v_{t+1}}^{2}.$
@@ 3.2 Linear consumption function parameters
The moments of the distribution of consumption growth reflect the underlying distributions of income shocks, as well as the distribution of unobserved taste heterogeneity. In general, larger dispersion, skewness, or kurtosis in the distributions of income shocks imply larger dispersion, skewness, or kurtosis in the distribution of consumption growth. The link between them depends on the degree of insurance with respect to these shocks, so these moments provide restrictions that help identify the parameters of partial insurance.
BPP discuss identification in the linear function at length; we provide a summary here while we relegate some details to appendix B.2. Formally, the transmission parameter of permanent shocks is identified as
@@
in both cases, the variance of shocks is given by the earlier expressions. [^14] The variance of consumption taste heterogeneity is further given by $\sigma_{\xi_{t}}^{2}={\operatorname*{\mathbb{E}}}((\Delta c_{it})^{2})-(\phi_{t}^{(1)})^{2}\sigma_{\zeta_{t}}^{2}-(\psi_{t}^{(1)})^{2}\sigma_{v_{t}}^{2}$. Its skewness and kurtosis are similarly identified from higher moments of consumption growth. These statements reflect the intuition behind identification in BPP: the covariance between consumption and (appropriately defined) income growth captures the strength of the relation between the two series and, consequently, the pass-through of income shocks into consumption. [^15] An equivalent view is that the partial insurance parameters are identified from a regression of consumption growth on income growth, using permanent or future income as instruments for income growth.
@@ 3.3 Quadratic consumption function parameters
As per the quadratic specification in (5) and the properties of shocks, the second moment of consumption growth is given by
@@
where square matrix $\mathbf{A}$ depends exclusively on the second, third, and fourth moments of income shocks. [^16] The matrix of coefficients is given by $\mathbf{A}=\left(\begin{array}[]{lllll}\sigma_{\zeta_{t}}^{2}&\sigma_{v_{t}}^{2}&\gamma_{\zeta_{t}}&\gamma_{v_{t}}&0\\ \gamma_{\zeta_{t}}&\gamma_{v_{t}}&\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{v_{t-1}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t-1}}^{2})&2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ 0&-\sigma_{v_{t}}^{2}&0&-\gamma_{v_{t}}&0\\ 0&\gamma_{v_{t}}&\sigma_{\zeta_{t}}^{2}(\sigma_{\zeta_{t+1}}^{2}+\sigma_{v_{t}}^{2}+\sigma_{v_{t+1}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(\sigma_{\zeta_{t+1}}^{2}+\sigma_{v_{t+1}}^{2})&0\\ 0&-\gamma_{v_{t}}&-\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}&-\kappa_{v_{t}}&-\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ \end{array}\right).$ , [^17] The variance of consumption taste heterogeneity is identified from ${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{2})$. Its skewness and kurtosis enter the third and fourth moments of consumption growth. Those depend on higher than fourth-order moments of income, which we do not model, so they are generally not identified in the current setting. We establish in appendix B.3 that matrix $\mathbf{A}$ is full rank and invertible for any distribution of income shocks except Bernoulli (for which there is no empirical evidence anyway). [^18] If the distributions of both shocks are Bernoulli, then there exist some parameter configurations for which matrix $\mathbf{A}$ is not full rank. Assuming away Bernoulli distributions is thus an identifying assumption. Identification of the parameters does not require those shocks to be non-Gaussian; the matrix is invertible even if the distributions of shocks are normal and have the same variance between them (i.e., $\sigma_{\zeta_{t}}^{2}=\sigma_{v_{t}}^{2}$).
@@
In the nonlinear specification, tail shocks exhibit a different pass-through compared to average shocks, regardless of the specific form of underlying income risk. As a result, their transmission parameters cannot be identified with second-order income moments alone. Skewness and kurtosis now become essential for identification, even if the income distribution is Gaussian. This is intuitive: we want to identify the impact of tail shocks, so we must use information on said tails. In appendix B.3, we explore how varying levels of income skewness and kurtosis influence the transmission parameters.
@@
Contrary to income, no higher-order moments of consumption are needed; the identifying moments in (8) are covariances between consumption growth and various powers of income growth. This is an advantage over quantile methods that require observing the entire consumption distribution, e.g., ABB, because, while we may estimate higher moments of consumption growth in the PSID, there is yet no established population benchmark to assess them against. This is in contrast to higher income moments, for which Guvenen et al. (2021) is a natural benchmark given their use of administrative population data.
@@
This fails to identify the true pass-through of permanent shocks in the presence of nonlinear transmission, i.e., expression (6). Moreover, this workhorse covariance is biased for $\phi_{t}^{(1)}$ if permanent shocks exhibit skewness. If $\gamma_{\zeta_{t}}<0$ and $\phi_{t}^{(2)}<0$, as our empirical estimates suggest, linearity biases $\phi_{t}^{(1)}$ upwards. Similarly, BPP’s identifying condition for $\psi_{t}^{(1)}$ becomes
@@
Our estimation procedure, which we describe in the next section, is linked transparently to the above identification argument. Effectively, we bring (8) to the data, along with some over-identifying restrictions, using moments estimation.